OpenStudy (anonymous):

How would I find the least common denominator of these fractions (posted below)...

OpenStudy (anonymous):

\[\frac{ \sqrt[3]{x} }{ 2\sqrt{x} } + \frac{ \sqrt{x}+3 }{ 3\sqrt[3]{x} }\]

OpenStudy (anonymous):

yaaa.. no

OpenStudy (anonymous):

lol That's exactly what I thought.

OpenStudy (anonymous):

at first i thought you be like 1/18 and 1/4 or something like that and I be like I know EXACTLY how to do that :) but ya... no...

OpenStudy (anonymous):

srry xD

OpenStudy (anonymous):

@triciaal @geny55

OpenStudy (anonymous):

See that's my problem, and I even know the answer (thanks to calculators). Problem is I have to show work, and I can't do that since I have no idea how to. But yeah. Don't worry about it. :)

OpenStudy (anonymous):

@pooja195 @perl

OpenStudy (anonymous):

@Ashleyisakitty @west_coast

OpenStudy (anonymous):

i can help

OpenStudy (anonymous):

It would be appreciated. :)

OpenStudy (anonymous):

ok, paste both fractions in separate comments so i can see them clearer

OpenStudy (anonymous):